# Exponential quantum advantages in learning quantum observables from classical data

### v1

@@ 3.1 Generalizations: quantum advantages for fixed inputs and for ground state problems
The “flipped case” As a final remark, it is interesting to observe that if we consider the case of fixed input state, which could be a fixed initial state $\rho_{0}(\bm{x})=\ket{\bm{x}}\bra{\bm{x}}$ | in the time evolution scenario modeled by the concept class in Eq. (1) or a single ground state $\rho_{H}(\bm{x})$ of a fixed Hamiltonian $H(\bm{x})$ in the ground state problem, then the learning problem becomes trivially classically easy. Such learning scenario is formally equivalent to considering a “flip concept” of the ones defined above. Let us for example consider the task of learning an unknown quantum process from many measurements, in this case the learning problem is still modeled by the concept class $\mathcal{F}^{H,O}_{\mathrm{evolved}}$ of Eq. (1) with the distinction that now the role of $x$ and ${\bm{\alpha}}$ are switched. Namely, the concepts are defined as $f^{\bm{x}}({\bm{\alpha}})$ so that the their expressions remain the same of $f^{\bm{\alpha}}(\bm{x})$. The difference lies in the labeling, where $x$ now denotes the concept while ${\bm{\alpha}}$ represents the input vectors. As $\bm{x}$ is constant for an instance of the learning problem, specified by the concept that generates the data, the training samples are measurements of the same quantum state $\rho_{H(\bm{x})}$ with different observables corresponding to different ${\bm{\alpha}}$. Since $O({\bm{\alpha}})=\sum_{j}{\bm{\alpha}}_{j}P_{j}$, we can make use of data to solve the linear system and obtain the expectation values of each local Pauli string $\text{Tr}[\rho_{H_{\mathrm{hard}}}(\bm{x})P_{i}]$. It then becomes easy to extrapolate the value of $\text{Tr}[\rho_{H_{\mathrm{hard}}}(\bm{x})O({\bm{\alpha}})]$ for every new ${\bm{\alpha}}$.
3.2 Generalization of the quantum learning mechanism to quantum kernels
@@
In [LAT21], considerable effort was invested to construct a task with provable learning speed-up where the quantum learning model is somewhat generic (related to natural QML models studied in literature), while still being capable of learning a classically unlearnable task. In our construction, it is possible to view the entire learning process as a quantum kernel setting, similarly to the approach in [YIM23]. However, in standard kernel approaches, especially those stemming from support vector machines, the optimization process is solved in the dual formulation where the hypothesis function is expressed as a linear combination of kernel functions. In contrast, the LASSO optimization employed here solves the optimization in the primal form with a constrain on the $\ell_{1}$ norm. This is an issue making our learner not technically a quantum kernel. While the LASSO formulation does not straightforwardly convert into a kernel method, one could attempt to address the regression problem outlined in Equation (6) using an alternative optimization approach that supports a kernel solution, such as kernel ridge regression. In this case the optimization is done for vectors with bounded $\ell_{2}$ norms, nevertheless there exist bounds on the generalization performance of such procedures as well [MRT18].
4 Generalization to observables parametrized by unitaries
@@
We have demonstrated the existence of a quantum advantage for the learning problem of predicting $k$-local observables from time evolved states and from ground states. Critically, we considered observables of the type $O({\bm{\alpha}})=\sum_{i}{\bm{\alpha}}_{i}P_{i}$ and we exploited the linear structure of $O({\bm{\alpha}})$ to ensure quantum learnability through LASSO regression. It is however natural to ask if quantum learnability can be achieved for other types of observables, while maintaining the classical hardness. In this section we consider the far more general case where the unknown observable is parameterized through a unitary matrix, i.e. $O({\bm{\alpha}})=W({\bm{\alpha}})OW^{\dagger}({\bm{\alpha}})$ where $O$ is an hermitian matrix. Our findings in this scenario will be of two kinds. First we show as a general result that for every method which learns a unitary $W$ given query access on a known distribution of input quantum states there exists a learning problem which exhibits a classical-quantum advantage. Then we concretize the general result by presenting a constructed example of a learning problem defined by a class of unitary-parameterized observables, showcasing a provable speed-up. Before stating our findings, let us properly introduce the learning problem under consideration. Imagine the scenario where the experimentalist once more measures observables on evolved quantum states. However, in this case, they lack control over the entire evolution, with a portion of it remaining unknown. Specifically, the experimentalist will receive measurements on input-dependent states |$\ket{\psi(\bm{x})}=W({\bm{\alpha}})U(\bm{x})\ket{0}$ for an unknown fixed ${\bm{\alpha}}$. Their objective remains predicting expectation values on states |$\ket{\psi(\bm{x}^{\prime})}$ for new inputs $\bm{x}^{\prime}$. Concretely such scenario corresponds to the following concept class:
$\mathcal{M}_{U,W,O}$ =$=\{f^{\bm{\alpha}}(\bm{x})\in\mathbb{R}\;\;|\;\;{\bm{\alpha}}\in[-1,1]\}$ (9)
with $\text{with }\;\;f^{\bm{\alpha}}(\bm{x}):\;x$ $\in\mathcal{X}\subseteq\{0,1\}^{n}\rightarrow\text{Tr}[\rho_{U}(\bm{x})O({\bm{\alpha}})]\in\mathbb{R}\;\;\&\;\;O({\bm{\alpha}})=W({\bm{\alpha}})OW^{\dagger}({\bm{\alpha}}).$ (10)
@@
where $\rho_{U}(\bm{x})=U(\bm{x})\ket{0}\bra{0}U^{\dagger}(\bm{x})$ and $O$ is a hermitian matrix. Given as training data $\mathcal{T}^{\bm{\alpha}}=\{(\bm{x}_{\ell},y_{\ell})\}_{\ell=1}^{N}$ with $\mathbb{E}[y_{\ell}]=\text{Tr}[\rho_{U}(\bm{x}_{\ell})O({\bm{\alpha}})]$, the goal of the learning algorithm is again to satisfy the learning condition of Def. 1. We are now ready to state our main general result of this Section:
Theorem 3 (Learning advantage for unitarily-parametrized observables).
@@
In other words, Theorem 3 guarantees that whenever there is an efficient method to learn a unitary from query access on a set of arbitrary input states and observables it is also possible to construct a learning problem of the kind Eq. (9) which exhibits a learning speed-up. We will more mathematically formalize the statement of Theorem 3 and provide a rigorous proof in the Appendix. The nontrivial part of our result lies in the fact that in the majority of the works present in the literature, which provide efficient algorithms to learn a unitary, the set of required probe states {|$\{\ket{\psi_{\ell}}\}_{\ell}$ is restricted to specific quantum states, often classically describable (e.g. stabilizer states). On such a set of states, our hardness results from the previous section can not directly be applied as a classical learning algorithm could simply prepare those states as the quantum algorithm would do.
The detailed proof of Theorem 3 can be found in the Appendix, here we outline the main ideas.
@@
The main idea of the proof is to construct a concept class of the kind of Eq.(9) which can be learned by a quantum algorithm using the learning algorithm $\mathcal{A}_{W}$, while maintaining classical harndess. Consider the circuit in Figure 1. Let $U(\bm{x})$ be a unitary that, depending on the first bit $x_{1}$ of the input $\bm{x}\in\{0,1\}^{n}$, prepares the |$\ket{\psi^{n_{S}}_{U}(\bm{x})}$ on the $n_{S}$ qubit register in two different ways. If $x_{1}=0$, then |$\ket{\psi^{n_{S}}_{U}(\bm{x})}$ is exactly one of the states |$\ket{\psi_{\ell}}\in S$, labeled by the final $n_{S}$ bits $\bm{x_{S}}\in\{0,1\}^{n_{S}}$ of the input bitstring $\bm{x}$, i.e. $\bm{x_{S}}=x_{n-n_{S}}...x_{n}$. If $x_{1}=1$, then $U(\bm{x})$ prepares the state |$\ket{\psi^{n_{S}}_{U}(\bm{x})}$ as the result of a BQP-complete computation. Regarding the observable, we define a controlled unitary $V_{A}$, controlled by the first $1+n_{Q}$ qubits register, such that when $x_{1}=0$ $V_{A}$ acts on the target $n_{S}$ qubit register by rotating the $n_{S}$ qubit measurement operator $O$ into one of the $Q_{m}$ in the set $Q$ required by the learning algorithm. The description of which $Q_{m}$ the unitary $V_{A}$ implements is contained in the $n_{Q}$ bitstring $\bm{x_{Q}}\in\{0,1\}^{n_{Q}}$ with $\bm{x_{Q}}=x_{2}x_{3}...x_{n_{Q}+1}$. When $x_{1}=1$, $V_{A}$ acts as the identity matrix on the $n_{S}$ qubit register. Suppose now the input bits are sampled from the following distribution $\mathcal{D}$ on $\bm{x}\in\{0,1\}^{n}$:
@@
  • If $x_{1}=1$ then the $n_{S}$ bits in $\bm{x_{S}}$ are sampled according to a hard distribution which makes the BQP computation of $U(\bm{x})$ classically hard on average. As discussed before, such distribution always exists for BQP-complete problems assuming the existence of an input distribution $\mathrm{BQP}\not\subset\mathrm{HeurP}^{\mathcal{D}^{\prime}}/\mathrm{poly}$.
Taking $W(\bm{0})=I^{\otimes n_{S}}$ for $\bm{\alpha}=\bm{0}$ and $O=I\otimes I^{\otimes n_{Q}}\otimes Z\otimes I^{n_{S}-1}$, it is clear that the concept $f^{0}=\text{Tr}[\rho_{U}(\mathbf{x})V_{A}OV^{\dagger}_{A}]$ cannot be learned by any classical algorithm. In particular, by the result of our Lemma 1 any classical algorithm can not meet the learning condition of Def.(1) on more than half of the inputs $\bm{x}$ coming from $\mathcal{D}$, namely it is guaranteed to fail for the inputs for which $x_{1}=0$. To prove quantum learnability we notice that for every $\bm{\alpha}$ the dataset $\mathcal{T}^{{\bm{\alpha}}}$ associated to the concept class $\mathcal{M}_{U,W,O}$ contains exactly the pairs of state and measurement outcomes needed by the learning algorithm $\mathcal{A}_{W}$ in order to learn the unitary $W(\bm{\alpha})$. Namely, half of the training samples, characterized by $x_{1}=0$ in their input $\bm{x}$, allows the algorithm to recover the unknown $W(\bm{\alpha})$. Therefore the quantum algorithm is able to learn the unknown observable $O({\bm{\alpha}})$ and evaluate it on each quantum state associated to every input $\bm{x}\in\{0,1\}^{n}$. ∎
As a corollary of Theorem 3, we construct a concrete example of learning advantage for observables parametrized by shallow unitaries by exploiting the recent results on learning shallow unitaries in [HLB+24].
@@
There exists a family of parametrized unitaries $\{U(\bm{x})\}_{x}$ and parametrized shallow circuits $\{W({\bm{\alpha}})\}_{\bm{\alpha}}$, a measurement $O$ and a distribution $\mathcal{D}$ over $x\in\{0,1\}^{n}$ such that the concept class $\mathcal{M}_{U,W,O}$ is not classically learnable with respect to the input distribution $\mathcal{D}$. However, there exists a quantum algorithm which learns $\mathcal{M}_{U,W,O}$ on the input distribution $\mathcal{D}$.
The full proof of Corollary 1 can be found in the Appendix.
{quantikz}
@@
[…]
4.1 Relationship to Hamiltonian learning
In order to further elucidate the lines between settings with and without classical/quantum learning advantages, we can establish a connection between the task of learning Hamiltonians from time-evolved quantum states and learning unitarily-parametrized observables. The setting of the Hamiltonian learning problem is the following. Given an unknown local Hamiltonian in the form of $H(\bm{\lambda})=\sum_{i}\lambda_{i}P_{i}$, the objective of the Hamiltonian learning procedure is to recover $\bm{\lambda}$. In the version of the problem we consider, we are given a black box which implements the time evolution under the unknown Hamiltonian $H(\bm{\lambda})$ on any arbitrary quantum state. The black box action on arbitrary inputs $\rho$ is given by

### v2

@@ IV.3 Generalizations: quantum advantages for fixed inputs and for ground state problems
The “flipped case” As a final remark, it is interesting to observe that if we consider the case of fixed input state, which could be a fixed initial state $\rho_{0}(\bm{x})=\ket{\bm{x}}\bra{\bm{x}}$ | in the time-evolution scenario modeled by the concept class in Eq. (3) or a single ground state $\rho_{H}(\bm{x})$ of a fixed Hamiltonian $H(\bm{x})$ in the ground state problem, then the learning problem becomes trivially classically easy. Such learning scenario is formally equivalent to considering a “flipped concept” of the ones defined above. Let us for example consider the task of learning an unknown quantum process from many measurements, in this case the learning problem is still modeled by the concept class $\mathcal{F}^{H,O}_{\mathrm{evolved}}$ of Eq. (3) with the difference that now the role of $\bm{x}$ and ${\bm{\alpha}}$ are switched. Namely, the concepts are defined as $f^{\bm{x}}({\bm{\alpha}})$ so that the their expressions remain the same of $f^{\bm{\alpha}}(\bm{x})$. The difference lies in the labeling, where $\bm{x}$ now denotes the concept while ${\bm{\alpha}}$ represents the input vectors. As $\bm{x}$ is constant for an instance of the learning problem, specified by the concept that generates the data, the training samples are measurements of the same quantum state $\rho_{H(\bm{x})}$ with different observables corresponding to different ${\bm{\alpha}}$. Since $O({\bm{\alpha}})=\sum_{j}{\bm{\alpha}}_{j}P_{j}$, we can make use of data to solve the linear system and obtain the expectation values of each local Pauli string $\text{Tr}[\rho_{H_{\mathrm{hard}}}(\bm{x})P_{i}]$. It then becomes easy to extrapolate the value of $\text{Tr}[\rho_{H_{\mathrm{hard}}}(\bm{x})O({\bm{\alpha}})]$ for every new ${\bm{\alpha}}$.
IV.4 Generalization of the quantum learning mechanism to quantum kernels
@@
In [3], considerable effort was invested to construct a task with provable learning speed-up where the quantum learning model is somewhat generic (related to natural QML models studied in literature), while still being capable of learning a classically unlearnable task. In our construction, it is possible to view the entire learning process as a quantum kernel setting, similarly to the approach in [14]. However, in standard kernel approaches, especially those stemming from support vector machines, the optimization process is solved in the dual formulation where the hypothesis function is expressed as a linear combination of kernel functions. In contrast, the LASSO optimization employed here solves the optimization in the primal form with a constrain on the $\ell_{1}$ norm. This is an issue making our learner not technically a quantum kernel. While the LASSO formulation does not straightforwardly convert into a kernel method, one could attempt to address the regression problem outlined in Equation (7) using an alternative optimization approach that supports a kernel solution, such as kernel ridge regression. In this case the optimization is done for vectors with bounded $\ell_{2}$ norms, nevertheless there exist bounds on the generalization performance of such procedures as well [17].
V Generalization to observables parametrized by unitaries
@@
We have demonstrated the existence of a quantum advantage for the learning problem of predicting $k$-local observables from time evolved states and from ground states. Critically, we considered observables of the type $O({\bm{\alpha}})=\sum_{i}{\bm{\alpha}}_{i}P_{i}$ and we exploited the linear structure of $O({\bm{\alpha}})$ to ensure quantum learnability through LASSO regression. It is however natural to ask if quantum learnability can be achieved for other types of observables, while maintaining the classical hardness. In this section we consider the far more general case where the unknown observable is parameterized through a unitary matrix, i.e. $O({\bm{\alpha}})=W({\bm{\alpha}})OW^{\dagger}({\bm{\alpha}})$ where $O$ is an hermitian matrix. Our findings in this scenario will be of two kinds. First we show as a general result that for every method which learns a unitary $W$ given query access on a known distribution of input quantum states there exists a learning problem which exhibits a classical-quantum advantage. Then we concretize the general result by presenting a constructed example of a learning problem defined by a class of unitary-parameterized observables, showcasing a provable speed-up. Before stating our findings, let us properly introduce the learning problem under consideration. Imagine a scenario where observables are measured on evolved quantum states, but there is no control over the entire evolution, with a portion of it remaining unknown. Specifically, measurements are taken on input-dependent states |$\ket{\psi(\bm{x})}=W({\bm{\alpha}})U(\bm{x})\ket{0}$ for an unknown fixed ${\bm{\alpha}}$. The goal is to predict expectation values on states |$\ket{\psi(\bm{x}^{\prime})}$ for new inputs $\bm{x}^{\prime}$. Concretely such scenario corresponds to the following concept class:
$\mathcal{M}_{U,W,O}=\{f^{\bm{\alpha}}(\bm{x})\in\mathbb{R}\;\;|\;\;{\bm{\alpha}}\in[-1,1]^{m}\}$ (9)
with: $\;\;f^{\bm{\alpha}}(\bm{x}):\;x\in\mathcal{X}\subseteq\{0,1\}^{n}\rightarrow\text{Tr}[\rho_{U}(\bm{x})O({\bm{\alpha}})]\in\mathbb{R}$
$O({\bm{\alpha}})=W({\bm{\alpha}})OW^{\dagger}({\bm{\alpha}}).$
@@
where $\rho_{U}(\bm{x})=U(\bm{x})\ket{0}\bra{0}U^{\dagger}(\bm{x})$ and $O$ is a hermitian matrix. Given as training data $\mathcal{T}^{\bm{\alpha}}=\{(\bm{x}_{\ell},y_{\ell})\}_{\ell=1}^{N}$ with $\mathbb{E}[y_{\ell}]=\text{Tr}[\rho_{U}(\bm{x}_{\ell})O({\bm{\alpha}})]$, the goal of the learning algorithm is again to satisfy the learning condition of Def. 4. We are now ready to state our main general result of this Section:
Theorem 5 (Informal version).
@@
In other words, Theorem 5 guarantees that whenever there is an efficient method to learn a unitary from query access on a set of arbitrary input states and observables it is also possible to construct a learning problem described by the concept class in Eq. (9) which exhibits a learning speed-up. We will more mathematically formalize the statement of Theorem 5 and provide a rigorous proof in the Appendix D. The nontrivial part of our result lies in the fact that in the majority of the works present in the literature, which provide efficient algorithms to learn a unitary, the set of required probe states {|$\{\ket{\psi_{\ell}}\}_{\ell}$ is restricted to specific quantum states, often classically describable (e.g. stabilizer states). On such a set of states, our hardness results from the previous section can not directly be applied as a classical learning algorithm could simply prepare those states as the quantum algorithm would do.
The detailed proof of Theorem 5 can be found in the Appendix D, here we outline the main ideas.
@@
The main idea of the proof is to construct a concept class of the kind of Eq.(9) which can be learned by a quantum algorithm using the learning algorithm $\mathcal{A}_{W}$, while maintaining classical harndess. Consider the circuit in Figure 1. Let $U(\bm{x})$ be a unitary that, depending on the first bit $x_{1}$ of the input $\bm{x}\in\{0,1\}^{n}$, prepares the |$\ket{\psi^{n_{S}}_{U}(\bm{x})}$ on the $n_{S}$ qubit register in two different ways. If $x_{1}=0$, then |$\ket{\psi^{n_{S}}_{U}(\bm{x})}$ is exactly one of the states |$\ket{\psi_{\ell}}\in S$, labeled by the final $n_{S}$ bits $\bm{x_{S}}\in\{0,1\}^{n_{S}}$ of the input bitstring $\bm{x}$, i.e. $\bm{x_{S}}=x_{n-n_{S}}...x_{n}$. If $x_{1}=1$, then $U(\bm{x})$ prepares the state |$\ket{\psi^{n_{S}}_{U}(\bm{x})}$ as the result of a $\operatorname{\mathsf{BQP}}$-complete computation. Regarding the observable, we define a controlled unitary $V_{A}$, controlled by the first $1+n_{Q}$ qubits register, such that when $x_{1}=0$ $V_{A}$ acts on the target $n_{S}$ qubit register by rotating the $n_{S}$ qubit measurement operator $O$ into one of the $Q_{m}$ in the set $Q$ required by the learning algorithm. The description of which $Q_{m}$ the unitary $V_{A}$ implements is contained in the $n_{Q}$ bitstring $\bm{x_{Q}}\in\{0,1\}^{n_{Q}}$ with $\bm{x_{Q}}=x_{2}x_{3}...x_{n_{Q}+1}$. When $x_{1}=1$, $V_{A}$ acts as the identity matrix on the $n_{S}$ qubit register. Suppose now the input bits are sampled from an arbitrary distribution $\mathcal{D}_{i}\in\{\mathcal{D}_{i}\}_{i}$ of the following kind:
@@
  • If $x_{1}=1$ then the $n_{S}$ bits in $\bm{x_{S}}$ follow any possible distribution over $\bm{x_{S}}\in\{0,1\}^{n_{S}}$. For each $i$, the distribution over $\bm{x_{S}}\in\{0,1\}^{n_{S}}$ defines the specific overall distribution $\mathcal{D}_{i}$.
Taking $W(\bm{0})=I^{\otimes n_{S}}$ for $\bm{\alpha}=\bm{0}$ and $O=I\otimes I^{\otimes n_{Q}}\otimes Z\otimes I^{n_{S}-1}$, it is clear that the concept $f^{0}=\text{Tr}[\rho_{U}(\mathbf{x})V_{A}OV^{\dagger}_{A}]$ cannot be learned by any classical algorithm. In particular, by the result of our Theorem 2 any classical algorithm can not meet the learning condition of Def. 4 on more than half of the inputs $\bm{x}$ coming from any distribution $\mathcal{D}_{i}$, namely it is guaranteed to fail for the inputs for which $x_{1}=1$. To prove quantum learnability we notice that for every $\bm{\alpha}$ the dataset $\mathcal{T}^{{\bm{\alpha}}}$ associated to the concept class $\mathcal{M}_{U,W,O}$ contains exactly the pairs of state and measurement outcomes needed by the learning algorithm $\mathcal{A}_{W}$ in order to learn the unitary $W(\bm{\alpha})$. Namely, half of the training samples, characterized by $x_{1}=0$ in their input $\bm{x}$, allows the algorithm to recover the unknown $W(\bm{\alpha})$. Therefore the quantum algorithm is able to learn the unknown observable $O({\bm{\alpha}})$ and evaluate it on each quantum state associated to every input $\bm{x}\in\{0,1\}^{n}$. ∎
V.1 Learning advantages for shallow unitaries
Theorem 5 presents a method for constructing learning problems that demonstrate a provable quantum advantage in learning unitarily parametrized observables of the form $O({\bm{\alpha}})=W({\bm{\alpha}})OW^{\dagger}({\bm{\alpha}})$. This result crucially relies on the existence of a learning scheme capable of recovering the unitary $W({\bm{\alpha}})$ based on output measurements performed on an arbitrary set of input states. As a concrete example where such a learning scheme is known to exist, and thus our result applies, we consider the case where the family of unitaries $\{W(\bm{\alpha})\}_{\bm{\alpha}}$ consists of shallow circuits. In this case, recent results in the literature [34] provide a procedure to learn few bodies observables from measurement of them on the single qubit stabilizer states. As shallow circuits preserve the locality of the observable measured, we can directly apply the result in [34] to construct a quantum learning algorithm for the concept class in Eq. (9) where the concepts are related to observables $O({\bm{\alpha}})=W({\bm{\alpha}})OW^{\dagger}({\bm{\alpha}})$, with $W({\bm{\alpha}})$ a shallow unitary. In particular, in the Appendix D we prove the following corollary of Theorem 5.
@@ V.1 Learning advantages for shallow unitaries
There exists a family of parametrized unitaries $\{U(\bm{x})\}_{x}$ and parametrized shallow circuits $\{W({\bm{\alpha}})\}_{\bm{\alpha}}$, a measurement $O$ and a set of distributions $\{\mathcal{D}_{i}\}_{i}$ over $\bm{x}\in\{0,1\}^{n}$ such that the concept class $\mathcal{M}_{U,W,O}$ is not classically learnable with respect to the set of input distributions $\{\mathcal{D}_{i}\}_{i}$, unless $\operatorname{\mathsf{BQP}}\subseteq\mathsf{P/poly}$. However, there exists a quantum algorithm which learns $\mathcal{M}_{U,W,O}$ on the input distributions $\{\mathcal{D}_{i}\}_{i}$.
Proof sketch.
The complete proof of Corollary 1 is provided in Appendix D. The proof follows the same approach as in Theorem 5, namely constructing a family of input distributions $\{\mathcal{D}_{i}\}_{i}$ that allows a quantum algorithm to recover the target unitary $W(\bm{\alpha})$, while ensuring that the learning problem remains intractable for any classical algorithm. Specifically, this is achieved by designing a model that prepares stabilizer states for half of the inputs and performs a $\operatorname{\mathsf{BQP}}$-hard computation for the other half. The quantum learning algorithm can then utilize the training data containing stabilizer states to learn the unitary $W(\bm{\alpha})$ by applying the procedure in [34]. Importantly, it will also be able to evaluate the learned $O(\bm{\alpha})$ on the other half of input data. ∎
@@
V.2 Relationship to Hamiltonian learning
In order to further elucidate the lines between settings with and without classical/quantum learning advantages, we can establish a connection between the task of learning Hamiltonians from time-evolved quantum states and learning unitarily-parametrized observables. The setting of the Hamiltonian learning problem is the following. Given an unknown local Hamiltonian in the form of $H(\bm{\lambda})=\sum_{i}\lambda_{i}P_{i}$, the objective of the Hamiltonian learning procedure is to recover $\bm{\lambda}$. In the version of the problem we consider, we are given a black box which implements the time-evolution under the unknown Hamiltonian $H(\bm{\lambda})$ on any arbitrary quantum state. The black box action on arbitrary inputs $\rho$ is given by